HELP Please!!!!! How many values do they want and what would they be?
What values for x must be excluded in the following fraction?

(x+5)/-1
THANKS!! :o)

Found 2 solutions by stanbon, tutorcecilia:
Answer by stanbon(75887)

x may be any Real Number; no Real Number is excluded.
Cheers,
Stan H.

Answer by tutorcecilia(2152)

(x+5)/-1
In this case, no real numbers need to be excluded.
The question is asking, what real number(s) will make this an untrue statement. The answer is "none" because all real numbers can be included. There are no exceptions.
For example, pick any real number and substitute it for "x", you will always get a real number as the answer:
(x+5)/-1
Let x=0
Than
(0+5)/-1=5/-1=-5
.
Try x=-2
Than (-2+5)/-1=3/-1=-3
.
Try x= -5
(-5+5)/-1=0/-1=0. It is OK to have zero in the numerator, but not in the denominator (an undefined condition)
.
But, if you get a problem such as:
-1/(x+5), than x can equal all real numbers EXCEPT FOR (-5). Why? Because:
-1/(-5+5)=-1/0=undefined. You cannot divide by zero.